l Review 1 Constanits1 Parodic Table Part A Find the acceleration of the two masses. Consider two masoes ng and connected by a thin sting. Assume the folowing values: m1-7.9kg and ma2-1.0 kg Ignore triction and the mass of the sting Value Units Submit Request Answer Part B- Changing the value of mass mi while keeping mass m2 constant would change the acceleration of the two-mass system. What should be the value of mass mj to get the largest possible value of acceleration of the two masses? mValue Units Submit Request Answer Part C What would be the value of that maximum acceleration? Value Units
Answer #1

A] Let the tension be T and acceleration be a, by force equation on two blocks,

m2g - T = m2a

T = m1a

adding both, m2g = (m1+m2)a

a = m2g/(m1+m2) = 1*9.8/[1+7.9]

= 1.1 m/s^2

B]   a = m2g/(m1+m2)

for largest a, m1 = 0 kg

C] a = g = 9.8 m/s^2